OpenStudy (anonymous):

How would i change this into slope intercept form? –x + 3y = 9

OpenStudy (anonymous):



OpenStudy (anonymous):

Slope-Intercept form: y = mx + b So solve for y: –x + 3y = 9 add x 3y = x + 9 divide by 3 y = 1/3x + 3

OpenStudy (anonymous):

When you change into slope-intercept form, you're just isolating "y" to one side of the equation.

OpenStudy (anonymous):

Thankyou very much!

OpenStudy (anonymous):

welcome (: